# The State and Fate of Summarization Datasets

### v1

@@ 4.5 Copyright Concerns Limit the Distribution of Datasets
Our survey shows that there is a recent trend away from publishing datasets in a public manner, instead making them available through reconstruction scripts from URLs (Grusky et al., 2018; Varab and Schluter, 2020; Kryściński et al., 2021), which may be due to growing legal concerns (Karamolegkou et al., 2023).
Reconstruction may be a temporary solution, which does not address the underlying legal issues. Moreover, this practice may lead to inconsistencies between local instances of the same dataset obtained at different times, due to changes in URL contents.
Recently, even datasets which were previously publicly available are being called into question. For example, the popular CNN/Daily Mail and XSUM are considered to be “legally questionable” Wang et al. (2022a). In a similar vein, the New York Times Corpus Sandhaus (2008), despite being published by the newspaper itself, was recently withdrawn and is no longer available. This may be another downside of the field’s over-reliance on the news domain.
@@ 5 Standardizing Summarization Datasets
Our survey surfaces a lack in terminology and reporting of the vast number of summarization datasets. This challenges both users of datasets, who may find it hard knowing whether there exists a dataset specifically tailored to their needs (e.g., in terms of language, domain, or shape), as well as researchers in the field, looking to identify current bottlenecks and directions for future research. Here we outline two contributions we make towards making datasets more standard and accessible.
Interactive web interface.
We will make available a web UI populated with our ontology of 133 summarization datasets, along with our annotations of the seven axes which characterize each dataset (See Figure 7). Users of this interface can search for a dataset by specifying the criteria they are looking for. We will welcome community contributions to add new datasets, or past datasets not already included in it. We hope that this interface can become a central hub, allowing users to easily find relevant datasets, without the need to manually sift through hundreds of papers.
Template for a summarization data card.
Specifically, we focus on four key elements to report: (1) sample information: shape, languages, domain; (2) annotation information, indicating how were the summaries created: e.g., via scraping or manual effort; (3) data quality assessment: we propose to report the level of abstraction via n-gram novelty, the Compression rate (Grusky et al., 2018), as well as human evaluation of the data, if it was performed; and (4) availability: indicating copyright licensing and download information.
6 Related Work
Several surveys have explored different aspects of automatic summarization Kryściński et al. (2019); Cajueiro et al. (2023); Sharma and Sharma (2022), yet their focus was more on models rather than datasets. To the best of our knowledge, none of the previous work has focused on the state of low-resource languages, nor looked at the dataset collection method and distribution. Dernoncourt et al. (2018) published a repository of available corpora for summarization. However, it included only a limited number of datasets (21), with just four that are not in English.
@@ 6 Related Work
A few papers focused on assessing the properties of popular datasets. Bommasani and Cardie (2020) conducted evaluation of 10 famous summarization datasets using novel metrics and Tejaswin et al. (2021) manually analyse 600 samples from three datasets.
@@ 8 Limitations
Second, we decided to omit from our ontology some axis. Specifically, we did not analysed data about sub-tasks of summarizations such as multi-modal, long-form or aspect based. We note that the available resources are still included in our platform and marked accordingly.
@@ Acknowledgments
We would like to thank Omer Kidron, Daria Lioubashevski and the members of the SLAB group at the Hebrew University for the productive discussion.

### v2

@@ 4.5 Copyright Concerns Limit the Distribution of Datasets
Our survey shows that there is a recent trend moving away from publishing datasets in a public manner, instead making them available through reconstruction from URLs (Grusky et al., 2018; Varab and Schluter, 2020; Kryściński et al., 2021), which may be due to growing legal concerns (Karamolegkou et al., 2023).
Reconstruction may be a temporary solution, which does not address the underlying legal issues. Moreover, it may lead to inconsistencies between local instances of the same dataset obtained at different times, due to changes in URL contents. This issue is sometimes mitigated by using snapshots.
Even datasets which were previously publicly available are being called into question. For example, the popular CNN/Daily Mail and XSUM are considered to be “legally questionable” Wang et al. (2022a). In a similar vein, the New York Times Corpus Sandhaus (2008), despite being published by the newspaper itself, was recently withdrawn and is no longer available. This may be another downside of the field’s over-reliance on the news domain.
@@ 5 Standardizing Summarization Datasets
Our survey surfaces a lack in terminology and reporting across the vast number of summarization datasets. This challenges researchers both in determining whether there exists a dataset specifically tailored to their needs (e.g., in terms of language, domain, or shape), and in identifying bottlenecks and directions for future research. Here we outline two contributions we make towards making datasets more standard and accessible.
Figure 7: Our interface allows researchers to determine if their specific needs are met by available resources. It is especially important for low resource languages.
Interactive web interface.
We make available a web UI populated with our ontology of 133 summarization datasets, along with our annotations of the seven axes that characterize each dataset (see Figure 7). Users of this interface can search for datasets by specifying the criteria they are looking for. We welcome community contributions to add new datasets, or past datasets not already included in it. We hope that this interface can allow users to easily find relevant datasets, without the need to manually sift through hundreds of papers.
Template for a summarization data card.
Specifically, we focus on four key elements to report: (1) sample information: shape, languages, domain; (2) annotation information, indicating how were the summaries created: e.g., via scraping or manual effort; (3) data quality assessment: we propose to report the level of abstraction via n-gram novelty, the Compression rate (Grusky et al., 2018), as well as human evaluation of the data, where applicable; and (4) availability: indicating copyright licensing and download information.
6 Related Work
@@
Several surveys have explored different aspects of automatic summarization Kryściński et al. (2019); Cajueiro et al. (2023); Sharma and Sharma (2022), yet they focus on models rather than datasets. To the best of our knowledge, none of the previous work has focused on the state of low-resource languages, nor looked at the dataset collection method and distribution. Dernoncourt et al. (2018) published a repository of available corpora for summarization. However, it included only a limited number of datasets (21), with just four that are not in English.
@@
A few papers focused on assessing the properties of popular summarization datasets. Bommasani and Cardie (2020) conducted an evaluation of 10 famous datasets using novel metrics and Tejaswin et al. (2021) manually analysed 600 samples from three datasets.
@@ 8 Limitations
Second, we decided to omit from our ontology some axes. Specifically, we did not analyse data about sub-tasks of summarizations such as multi-modal, long-form or aspect based. We note that the available resources are still included in our platform and marked accordingly.
@@ Acknowledgments
We would like to thank Omer Kidron, Daria Lioubashevski, the members of the SLAB group at the Hebrew University for the productive discussion, and the anonymous reviewers for their valuable feedback. This works was partially supported by research grants no. 2088, 2336, both from the Israeli Ministry of Science and Technology.